# LLMs Are In-Context Bandit Reinforcement Learners

### v2

@@ 1 Introduction
Large language models (LLMs) have been shown to exhibit in-context learning (ICL), a form of supervised learning that does not require parameter updates (Brown et al., 2020). ICL relies on including supervised input-output pairs in the LLM context (i.e., prompt), [^1] We use the terms prompt and context interchangeably. and it has proven effective with both few (Brown et al., 2020) and many (Bertsch et al., 2024; Agarwal et al., 2024) examples. We ask whether the ability to learn in-context extends to contextual bandit reinforcement learning (RL), i.e., whether language models can effectively perform in-context reinforcement learning (ICRL) with stateful single-step interaction episodes. This is a desirable property of LLMs, as they can learn new tasks interactively, in an online setting, at deployment time, without requiring demonstration data, but based on rewards on their outputs. This learning signal may be human-generated, automatically provided (i.e., a program successfully completes; Gehring et al., 2024), or even AI-generated (Zhang et al., 2024a).
@@
ICRL naturally combines ICL and reinforcement learning (RL). In contrast to ICL, as an RL process, ICRL does not rely on annotated labels or a fixed dataset. Instead of constructing the LLM context from supervised input-output pairs, the LLM context is constructed from triplets of an input, a model’s predicted output, and its reward. As more input examples are observed, the model has access to additional triplets, leading to an online and continual learning scenario, where model capabilities improve over time. These triplets are followed by a new input, for which the model must predict an output. Figure 1 illustrates bandit ICRL prompting.
@@
We study the ICRL capabilities of Llama 3.1 (Llama Team, 2024), Qwen2.5 (Qwen et al., 2024), and Gemini 1.5 Flash (Gemini Team, 2024) [^2] We also conduct early experiments on Phi-3.5-mini (Abdin et al., 2024), included in the appendix., using several many-label classification benchmarks, with a focus on contextual bandit RL (i.e., stateful single-step episodes).
@@
We find that LLMs demonstrate innate ICRL capabilities, but that two choices are critical for effective learning. First, a relatively high stochasticity level is needed to encourage exploration. Second, using only triples with positive rewards performs best, corresponding to how the classic REINFORCE algorithm (Williams, 1992) essentially ignores zero-reward examples. The latter choice creates a cosmetic similarity to ICL. However, ICRL remains fundamentally different: in ICRL the model must actively explore to find (i.e., generate) these positive triplet pairs, rather than having an expert annotator provide them.
@@
A recurring observation in our experiments is the relative instability of the process, as performance can suddenly dip significantly, before often quickly recovering. We propose a new method, Stochastic ICRL, to add stochasticity to the prompt construction by only sampling some of the past episodes observed in context, instead of increasing the temperature. This enhances exploration, stabilizes performance, and maintains relatively shorter contexts. Interestingly, this also allows the model to learn in the presence of negative signals as well, but still less effectively than without.
Finally, we observe the scaling trends of ICRL, by studying all Qwen2.5 modeling sizes between 0.5B and 72B. As expected, small models exhibit limited capabilities, while larger variants deliver the strongest performance.
Overall, we show that LLMs display ICRL abilities that lead to consistent and often substantial gains across all tasks we study, in both the naive and stochastic versions. For example, in the Banking77 (Casanueva et al., 2020) classification task, Qwen2.5-7B improves from 6.2% zero-shot accuracy to 72.2% through Stochastic ICRL. Our code, data, and experimental logs are available at https://lil-lab.github.io/icrl/.
@@ 2 In-context Reinforcement Learning
ICRL follows the reinforcement learning paradigm (Sutton & Barto, 2018), where models learn by reinforcing their own good behaviors and suppressing their own bad choices. Instead of providing models with correct demonstrations, the model generates an output given an input, then observe the outcome (i.e., reward) of its prediction. It learns from the reward signals, in an online learning setting, all within the context. In this study we focus on a contextual bandit RL scenario, a restricted variant of RL, where the length of each episode is one step.
@@ 2.1 Naive and Naive+ ICRL
0:  $\mathcal{D}$: Data distribution
0:  $\pi$: Language model policy
@@
Algorithm 1 describes the Naive approach, as well as Naive+, a variant that only considers examples with positive reward. Omitting lines 78 gives Naive ICRL, the most straightforward way to implement ICRL. At each time step $t$, the model observes a new example $x^{(t)}$, produces a prediction $\hat{y}^{(t)}$, and receives a reward $r^{(t)}$. Every such model interaction creates an episode, which is appended to the buffer $\mathcal{E}$. For each new interaction, we construct a context $C$ from prior episodes (line 3). In Naive, this context is simply all past episodes. Naive+ ICRL adds lines 78 and ignores negative-reward episodes, retaining only positive episodes in the buffer. In both variants, as the LLM’s context window becomes full, we drop older episodes to maintain a sliding window of the most recent interactions.
2.2 Stochastic ICRL
Stochastic ICRL utilizes model sensitivity to prompt composition as an avenue to increase exploration, instead of the increased temperature of Naive. Changes in prompt composition have been widely observed to lead to variance in LLM behavior, including through changes in the set of ICL examples (Zhang et al., 2022; Liu et al., 2022; Chen et al., 2023; Levy et al., 2023), seemingly meaningless stylistic changes (Sclar et al., 2024; Lu et al., 2022), or even interventions based on entropy in the embedding space (Rahn et al., 2024). Generally, this property of LLMs is not viewed positively. However, it introduces stochasticity into the ICRL process, and thereby introduces exploration.
Stochastic introduces context stochasticity by randomly choosing the subset of past episodes to include in the prompt each time the model observes a new input. As in Naive+ ICRL, Stochastic only uses episodes with positive reward, as we empirically observe it to lead to better results.
@@ 2.2 Stochastic ICRL
0:  $\mathcal{D}$: Data distribution
0:  $\pi$: Language model policy
@@
0:  $p_{\text{keep}}$: Prob. to keep examples in context
@@
Algorithm 2 describes Stochastic ICRL. For each input, we construct a new context (lines 39). We decide what past episodes to include in this context by sampling from a Bernoulli variable parameterized by $p_{\text{keep}}$ (lines 49). We sample independently for each past episode. This results in different implicit reasoning for each input, because each is done with a different context. When storing past episodes, we only include episodes with positive reward (lines 1315).
With small $p_{\text{keep}}$, Stochastic will encounter the issue of the LLM context window saturating much later than Naive. However, deploy ICRL for enough interactions, and the context window will saturate, even for models with the largest windows. Similar to naive, we downsample the context if it overflows the LLM context window. We do it by removing episodes from the sampled $C^{(t)}$ uniformly at random until the context fits the model’s context window.
@@ 3 Experimental Setup
We follow Bertsch et al.’s (2024) study of many-shot ICL in focusing on five classification problems: Banking77 (77 labels; Casanueva et al. (2020)), CLINC150 (150 labels; Larson et al. (2019), NLU (68 labels; Liu et al. (2021)), TREC (6 labels; Li & Roth (2002); Hovy et al. (2001)), and TREC-fine (50 labels; Li & Roth (2002); Hovy et al. (2001)). Because of the large output spaces (up to 150 labels in CLINC150), these tasks are particularly challenging for large language models, as empirically shown by Bertsch et al. (2024) and replicated in our zero-short results. The classification problem creates a contextual bandit scenario. The labels in each dataset are used to compute rewards, and are never shown to the model. Appendix C.3 provides additional information about the datasets.
@@
We simulate interactive binary rewards from perfect human actors interacting with the system. We do so by comparing the model outputs with the gold label of each input. This is a common practice in studies of the effect of rewards on learning processes (Gao et al., 2023; Lightman et al., 2024), for practical convenience. We deterministically transform the numerical rewards into a natural language format indicating if the model prediction is correct or not, which is more suitable for LLM reasoning. Appendix C.1 provides more details about our prompting.
@@ 4 Results and Analysis
Figure 3: Reward ablations. Test accuracies of Naive and Stochastic with different reward signals. Positive reward only is the best choice for both methods. With Naive, no other strategy facilitates learning. 3(b) in Appendix D details start and end accuracies.
@@
Both Naive+ and Stochastic effectively learn in all tasks and for both models, showing significant improvements over zero-shot. Naive+ and Stochastic improve over the zero-shot accuracies by between 28.6–74.4% for Llama, and 29.2–68.4% for Qwen. In general, accuracies approach the upper bound in many settings, demonstrating the strong bandit ICRL capabilities of Llama and Qwen. For both models, both approaches also demonstrates a monotonic growth in performance over time (Stochastic especially), suggesting that with more data performance would improve further. This is especially evident for the most challenging datasets, i.e., the ones with the most labels (i.e., Banking77, CLINC150, NLU), as they generally require more exploration to learn how to map the inputs to the large number of potential outputs.
[…]
@@
Overall, our ablations highlight two key points: (1) LLMs can learn online from their predictions only when a reward signal is present, and (2) LLMs exhibit inherent limitations in implicitly learning from mistakes (i.e., without explicit reasoning, as in Wei et al.’s (2022) Chain-of-Thought).
@@
Figure 4 summarizes our findings on Llama 8B, Qwen2.5 7B, and Gemini 1.5 Flash 8B. To contextualize the results, we include upper bound results from standard supervised ICL with as many gold demonstrations (with abstract labels) as the context can handle, which generally succeeds across tasks (except for a lower performance on TREC-fine). With just one exemplar per label, ICRL nearly reaches the upper bound for all tasks and models. We also stress that just including one gold demonstration per label is not always enough to reach good performance, and the online process is still important: for example, Llama with exemplars, when tested on CLINC150 with Stochastic, reaches non-negligeble accuracies only after 3k steps. In the absence of any exemplars, the ICRL process still manages to build informative contexts, though overall accuracy is not surprisingly lower. For instance, Qwen and Llama achieve higher than 45% accuracy on NLU and TREC with both Naive+ and Stochastic, indicating that the models learn and refine their output mappings over time. In contrast, Gemini excels when the correct mapping is provided but struggles significantly in a purely exploratory setting.
Given the domain effect observed with semantic labels, and the relatively low, even if significant results observed with abstract labels, the question arises if this learning is due to a domain effect. For example, Llama on CLINC150 improves from 0.8$\rightarrow$27.6during learning with Naive+, a significant, but modest improvement. We ran additional experiments to study the presence of learning with abstract labels, but without reward signals (i.e., to measure the domain effect). We see no improvement over zero-shot performance, confirming the reward signal is what drives learning. [^8] Because these experiments showed no learning effect at all, we are omitting them from our figures.
@@
An important differentiating factor between Stochastic and Naive+ is stability. The results so far (Figures 2, 4, and 5(a)) often show sudden, even if temporary dips in performance with Naive+. This instability is undesirable, because it means the performance of the model in interactions (i.e., with users) temporarily deteriorates significantly. In contrast, visually, the learning trends of Stochastic are more stable.
@@ 5 Related Work
Likely the closest work to ours is Krishnamurthy et al.’s (2024) study of whether LLMs can solve multi-armed bandit problems, a state-less simpler RL setting than the one we are focused on. We observe similar issues to their findings with the Naive approach. They present a set of negative results, and finally are able to elicit effective learning, but through a prompting strategy that cannot generalize beyond their very simple scenario. We address this challenge by showing the strong performances of both Naive+, which includes only positive outcomes and an increased sampling temperature, and Stochastic, which features stochasticity in the prompt construction. Concurrent to our work, Nie et al. (2024) studied the contextual bandit ICRL, similar to our study. They also propose a working solution, but take a very different approach by externally tracking learning statistics commonly used in the UCB bandit learning algorithm (Auer et al., 2002), and fine-tuning or prompting a model to leverage them. In contrast, the approaches we discuss do not rely on explicitly tracking statistics or fine-tuning. Instead, we are intered in innate abilities.
@@ 6 Discussion and Limitations
While our work provides a plethora of insights into ICRL behavior, much remain to be studied. We intentionally selected the contextual bandit setting and classification benchmarks and focused on binary rewards to simplify the experiments and evaluation in this early stage of studying ICRL. This is a limitation of this study, and leaves open the question of applicability to more complex RL problems, where rewards are more nuanced, or where interactions comprise multiple steps. For example, math and coding tasks often require multiple steps, but also introduce complex evaluation challenges. We believe our study enables future work to study these challenges, and that this is an important direction.
The limited capability we identified to implicitly reason about episodes with negative rewards in current LLMs also outlines an important research question for future work. Our work also leaves open questions about the use of computational resources. ICRL is relatively compute-intensive, especially after the learner observes many episodes. We propose Approximate ICRL in Appendix B.3 to reduce certain forms of computational overhead, and show how it allows to trade-off compute for robustness. Further reducing computational demands is an important direction for future work.

### v3

@@ 1 Introduction
Large language models (LLMs) have been shown to exhibit in-context learning (ICL), a form of supervised learning that does not require parameter updates (Brown et al., 2020). ICL relies on including supervised input-output pairs in the LLM context (i.e., prompt), [^1] We use the terms prompt and context interchangeably. and it has proven effective with few (Brown et al., 2020) and many (Bertsch et al., 2024; Agarwal et al., 2024) examples. We ask whether the ability to learn in-context extends to contextual bandit reinforcement learning (RL), i.e., whether language models can effectively perform in-context reinforcement learning (ICRL) with stateful single-step interaction episodes.
@@
ICRL naturally combines ICL and reinforcement learning (RL). In contrast to ICL, as an RL process, ICRL does not rely on annotated labels or a fixed dataset. Instead of constructing the LLM context from supervised input-output pairs, the LLM context is built from triplets of an input, a model’s predicted output, and its reward. As more input examples are observed, the model has access to additional triplets in context, leading to an online and continual learning scenario, where model capabilities improve over time. These triplets are followed by a new input, for which the model must predict an output. In this in-context framework, adding a past episode to the context corresponds, in standard fine-tuning RL settings, to using an episode at training time. Figure 1 illustrates bandit ICRL prompting.
ICRL is a desirable ability of LLMs. It allows learning new tasks interactively, in an online setting, at deployment time (without parameter updates), without requiring demonstration data. This learning signal may be human-generated, automatically provided (i.e., a program successfully completes; Gehring et al., 2024), or even AI-generated (Zhang et al., 2024a).
@@
We study the bandit ICRL capabilities of Llama 3.1 (Llama Team, 2024), Qwen2.5 (Qwen et al., 2024), and Gemini 1.5 Flash (Gemini Team, 2024). [^2] We also conduct early experiments on Phi-3.5-mini (Abdin et al., 2024), included in the appendix. Following existing bandit learning literature (Zhang et al., 2019; Bietti et al., 2021), we use many-label classification benchmarks to create contextual bandit RL scenarios, which simplify experimentation and evaluation, while focusing on the fundamental skills of exploration and learning from rewards.
@@
We find that LLMs demonstrate innate ICRL capabilities, but that two choices are critical for effective learning. First, a relatively high stochasticity level is needed to encourage exploration. Second, using only triples with positive rewards performs best.The latter choice creates a cosmetic similarity to ICL. However, ICRL remains fundamentally different: in ICRL the model must actively explore to find (i.e., generate) these positive triplets, rather than having an expert annotator provide them.
@@
A recurring observation in our experiments is the relative instability of the process, as performance can suddenly dip significantly, before often quickly recovering. We propose a new method, Stochastic ICRL, to add stochasticity to the prompt construction by only sampling some of the past episodes observed in context, instead of increasing the temperature. This enhances exploration, stabilizes performance, and maintains relatively shorter contexts. Interestingly, this also allows the model to learn in the presence of negative signals.We also study the scaling trends of ICRL, using all Qwen2.5 modeling sizes between 0.5B and 72B. There is a strong correlation between performance and model scale, but across all scales the relation between methods is maintained, with regard to both performance and stability.
Overall, we demonstrate that applying bandit ICRL consistently and significantly enhances the performance of LLMs. For example, in the Banking77 (Casanueva et al., 2020) classification task, Qwen2.5-7B improves from 6.2% zero-shot accuracy to 72.2% through Stochastic ICRL, without access to gold labels and without any updates of model parameters. These results suggest that LLM hold previously understudied capabilities for in-context learning, and lay the foundation for their further development and study in future work. Our code, data, and experimental logs are available at https://lil-lab.github.io/icrl/.
@@ 2 In-context Reinforcement Learning
ICRL follows the reinforcement learning paradigm (Sutton & Barto, 2018), where models learn by reinforcing their own good behaviors and suppressing their own bad choices. Instead of providing models with correct demonstrations, the model generates an output given an input, then observe the outcome (i.e., reward) of its prediction. It learns from the reward signals, in an online learning setting, all within the context (i.e., without parameter updates). In this study we focus on a contextual bandit RL scenario, a restricted variant of RL, where the length of each episode is one step.
@@ 2.1 Naive and Naive+ ICRL
[…]
@@
[…]
@@
[…]
@@
Algorithm 1 describes the Naive approach, as well as Naive+, a variant that only considers examples with positive reward. Omitting lines 78 gives Naive ICRL, the most straightforward way to implement ICRL. At each time step $t$, the model observes a new example $x^{(t)}$, produces a prediction $\hat{y}^{(t)}$, and receives a reward $r^{(t)}$. Every such model interaction creates an episode, which is appended to the buffer $\mathcal{E}$. For each new interaction, we construct a context $C$ from prior episodes (line 3). In Naive, this context is simply all past episodes. Naive+ ICRL adds lines 78 and ignores negative-reward episodes, retaining only positive episodes in the buffer. As the LLM’s context window fills, both variants maintain a sliding window by dropping older episodes.
2.2 Stochastic ICRL
Stochastic ICRL utilizes model sensitivity to prompt composition as an avenue to increase exploration, instead of the increased temperature of Naive. Changes in prompt composition have been widely observed to lead to variance in LLM behavior, including through changes in the set of ICL examples (Zhang et al., 2022; Liu et al., 2022; Chen et al., 2023; Levy et al., 2023), seemingly meaningless stylistic changes (Sclar et al., 2024; Lu et al., 2022), or even interventions based on entropy in the embedding space (Rahn et al., 2024). Generally, this property of LLMs is not viewed positively. However, it adds stochasticity to the ICRL process, which encourages exploration.
Stochastic introduces context stochasticity by randomly choosing the subset of past episodes to include in the prompt for each new input. Like Naive+ ICRL, it includes only positive-reward episodes, which improves results empirically.
Algorithm 2 describes Stochastic ICRL. For each input, we construct a new context (lines 39). We decide what past episodes to include in this context by sampling from a Bernoulli variable parameterized by $p_{\text{keep}}$ (lines 49). We sample independently for each past episode. This results in different implicit reasoning for each input, because each is done with a different context. As in Naive+, we only store episodes with positive reward (lines 1315).
With small $p_{\text{keep}}$, Stochastic will encounter the issue of the LLM context window saturating much later than Naive. However, deploy ICRL for enough interactions, and the context window will saturate, even for models with the largest windows.
Similar to naive, we downsample the context if it overflows the LLM context window. We do it by removing episodes from the sampled $C^{(t)}$ uniformly at random until the context fits the model’s context window.
@@ 3 Experimental Setup
We follow Bertsch et al.’s (2024) study of many-shot ICL in focusing on five classification problems: Banking77 (77 labels; Casanueva et al., 2020), CLINC150 (150 labels; Larson et al., 2019, NLU (68 labels; Liu et al., 2021), TREC (6 labels; Li & Roth, 2002; Hovy et al., 2001), and TREC-fine (50 labels; Li & Roth, 2002; Hovy et al., 2001). Because of the large output spaces (up to 150 labels in CLINC150), these tasks are particularly challenging for large language models, as empirically shown by Bertsch et al. (2024) and replicated in our zero-shot results. The classification problem creates a contextual bandit scenario (Zhang et al., 2019; Bietti et al., 2021). The labels in each dataset are used to compute rewards, and are never shown to the model. Appendix C.3 offers more details on the datasets.
@@
We simulate interactive binary rewards from perfect automatic verifiers or human actors interacting with the system. We do so by comparing the model outputs with the gold label of each input. This is a common practice in studies of the effect of rewards on learning processes (Gao et al., 2023; Lightman et al., 2024), for practical convenience. We deterministically transform the binary numerical rewards into a natural language format indicating if the model prediction is correct or not, which is more suitable for LLM reasoning. This formulation abstracts over challenges like exact numerical interpretation (i.e., of continuous rewards), while focusing on the fundamental skills of exploration and learning from rewards. Appendix C.1 elaborates on our prompting.
@@ 4 Results and Analysis
Both Naive+ and Stochastic effectively learn in all tasks and for both models, showing significant improvements over zero-shot. Naive+ and Stochastic improve over the zero-shot accuracies by between 28.6–74.4% for Llama, and 29.2–68.4% for Qwen. In general, accuracies approach the supervised performance in many settings, demonstrating the strong bandit ICRL capabilities of Llama and Qwen. Performance also grows monotonically over time, especially with Stochastic, suggesting further gains with more data. This trend is most evident in the most challenging datasets (Banking77, CLINC150, NLU), where high label counts demand more exploration to map inputs to outputs.
@@
Figure 3: Reward ablations. Test accuracies of Naive and Stochastic with different reward signals. Positive reward only is the best choice for both methods. With Naive, no other strategy facilitates learning. 12(d) in Appendix D details start and end accuracies.
@@
Overall, our ablations show that (a) LLMs can learn online from their predictions only when a reward signal is present, and (b) LLMs exhibit inherent limitations in implicitly learning from mistakes (i.e., without explicit reasoning, as in Wei et al.’s (2022) Chain-of-Thought).
[…]
@@
Figure 4 summarizes our findings on Llama 8B, Qwen2.5 7B, and Gemini 1.5 Flash 8B. To contextualize the results, we include upper bound results from standard supervised ICL with as many gold demonstrations (with abstract labels) as the context can handle, which generally succeeds across tasks (except for a lower performance on TREC-fine). With just one exemplar per label, ICRL nearly reaches the upper bound for all tasks and models. We also stress that just including one gold demonstration per label is not always enough to reach good performance, and the online process is still important: for example, Llama with exemplars, when tested on CLINC150 with Stochastic, reaches non-negligible accuracies only after 3k steps. In the absence of any exemplars, the ICRL process still manages to build informative contexts, though overall accuracy is not surprisingly lower. For instance, Qwen and Llama achieve higher than 45% accuracy on NLU and TREC with both Naive+ and Stochastic, indicating that the models learn and refine their output mappings over time. In contrast, Gemini excels when the correct mapping is provided but struggles significantly in a purely exploratory setting.
[…]
@@
An important differentiating factor between Stochastic and Naive+ is stability. The results so far (Figures 2, 4, and 5(a)) often show sudden, even if temporary dips in performance with Naive+. This instability is undesirable, because it means the performance of the model in interactions (i.e., with users) temporarily deteriorates significantly. In contrast, Stochastic’s learning trends are more stable.
@@ 5 Related Work
Likely the closest work to ours is Krishnamurthy et al.’s (2024) study of whether LLMs can solve multi-armed bandit problems, a state-less simpler RL setting than the one we are focused on. We observe similar issues to their findings with the Naive approach. They present a set of negative results, and finally are able to elicit effective learning, but through a prompting strategy that cannot generalize beyond their very simple scenario. We address this challenge by showing the strong performances of both Naive+, which includes only positive outcomes and an increased sampling temperature, and Stochastic, which features stochasticity in the prompt construction. Concurrent to our work, Nie et al. (2024) studied the contextual bandit ICRL, similar to our study. They also propose a working solution, but take a very different approach by externally tracking learning statistics commonly used in the UCB bandit learning algorithm (Auer et al., 2002), and fine-tuning or prompting a model to leverage them. In contrast, the approaches we discuss do not rely on explicitly tracking statistics or fine-tuning. Instead, we are interested in innate abilities.
@@ 6 Discussion and Limitations
Fundamentally, our work illustrates that exploration is the key ingredient necessary for ICRL behavior in LLMs. When exploration is combined with filtering of episodes with negative rewards, conventional ICL abilities (i.e., learning from demonstrations) bring about strong ICRL trends. Furthermore, exploration can be aided by introducing stochasticity in the prompt construction. The dependence of the learning trends on filtering out negative rewards leaves an important challenge for future work – how to elicit or train LLMs to reason effectively about negative episodes.
While our work provides a plethora of insights into ICRL behavior, much remains to be studied. We intentionally choose the contextual bandit setting using classification benchmarks following (Zhang et al., 2019; Bietti et al., 2021), and focus on binary rewards to simplify the experiments and evaluation in this early stage of studying ICRL. This formulation abstracts over challenges like exact numerical interpretation (i.e., of rewards), while focusing on the fundamental skills of exploration and learning from rewards. However, this limitation leaves open the question of applicability to more complex RL problems, where rewards are more nuanced, or where interactions comprise multiple steps. For example, math and coding tasks often require multiple steps, but also introduce complex evaluation challenges. We believe our study enables future work to study these challenges, and that this is an important direction.
Our work also leaves open questions about the use of computational resources. ICRL is relatively compute-intensive, especially after the learner observes many episodes. We propose Approximate ICRL in Appendix B.3 to reduce certain forms of computational overhead, and show how it allows to trade-off compute for robustness. Further reducing computational demands is an important direction for future work.